From GlueXWiki
Jump to: navigation, search

May 3, 2018 Drift Chamber meeting


  1. Instructions for Bluejeans meeting connection
  2. Meeting ID: 290664653

Headline text

  1. To join via a Web Browser, go to the page [1]


  1. CDC update
efficiency vs n, doca>0.35cm
efficiency vs n, doca>0.15cm
  1. FDC update
  2. NIM publications CDC pdf svn repository
  3. Other


Participants: Naomi(CMU), Nacer (GSI), Eugene, Simon, Alex A., Alex B., Beni, Cody, and Lubomir (JLab).


- Naomi showed plots demonstrating CDC gradually deteriorates its performance during Dec. 2017 - May 2018 running. The efficiency drops at big drift distances significantly; at the same time time-to-distance remains almost unchanged (Alex B.). The most probable explanation is gas contamination. It is strange that the hit efficiency exhibits 30deg phi variation. We discussed also other than gas explanations - some features in the tracking, electronics, etc. The plan is first to fix the gas leakage and than look with cosmics (also before fixing and disconnecting cables).

- Nacer finalized his studies of the dE/dx truncation (presentation attached). He simulated pions and protons and looked at the fraction of the misidentified protons for a fixed proton efficiency as a function of the momentum and angle. The minimum of this fraction is achieved for dE/dx truncation of 20-40%. A 20% truncation is used now in the analysis.

- Beni showed the effect of the W-foil on the CDC per-ring hit multiplicity. It is reduced significantly both for random and physics trigger and also similarly for inner and outer rings, indicating higher energy photons are contributing.


- The FDC current dropped by ~25% after insertion of the W-foil at the entrance pipe, to be compared to 5% drop when the foil was directly at the front of the AC. Thus, we are considering installing the W-foil in the tagger.

NIM publications

- Naomi went over the current version of the CDC paper. Needs contributions from other people (svn link attached above).


- Cody shared his plans for the summer and asked for an input what needs to be done. He has 10-20 spare modules that require some repair. ADCs can be fixed here, FPGAs by the company - require re-flow/re-ball $100-150/module - Eugene agreed. Work on the firmware - what should be the number of samples above threshold in the peak-finding algorithm (to avoid noise - to be discussed with Sergey), no need for high timing resolution after the first peak, implement the peak-finding procedure as it was used for the TRD studies, work on the self-triggering.